OpenStudy (anonymous):

Part 1: Use calories as the y-coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 : Write an equation, in slope intercept form, to represent this data. Part 3: How many calories will this person burn after 35 minutes of walking? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned. @Hyper @jim_thompson5910

OpenStudy (anonymous):

Oops, this was the first part: Assume a 115 pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking.

OpenStudy (hyper):

For the first part, do exactly as they tell you and sketch a graph with calories representing the y-axis and minutes representing the x - axis. From there, the second information you provided gives you the 2 ordered pairs or coordinates which they want you to use to find the slope. Plot these two points on the graph and the slope represents the; rise/run

OpenStudy (anonymous):

Which one is the ordered pairs?

OpenStudy (hyper):

They take the form of : ( minutes , calories burnt) which means that for each ordered pairs , the x coordinate will be the minutes and the y coordinate will be calories burnt. Ordered pair is just another term for a coordinate.

OpenStudy (anonymous):

Are you good at drawing lines?

OpenStudy (hyper):

not really...

OpenStudy (anonymous):

Okay, I'll make a second attempt at making a decent looking graph, hah.

OpenStudy (hyper):

Another hint: When t = 0 or x coordinate equals zero which is initially the time when the person walked, calories burnt or the y coordinate will also be zero.

OpenStudy (anonymous):

http://sketchtoy.com/32920037 Like this ?

OpenStudy (hyper):

The axis's are in the correct order. The question said that the x - coordinate which is horizontal represents the time and the y - coordinate which is the vertical represents the calories. Swap them around and apart from that, the information plotted and numbers is correct.

OpenStudy (anonymous):

Okay, one sec.

OpenStudy (anonymous):

http://sketchtoy.com/32921279 Like that?

OpenStudy (anonymous):

@jim_thompson5910 Can you help? This is my last question and I really want to get finished because I've been at this for almost 6 hours.

jimthompson5910 (jim_thompson5910):

the graph looks good to me, nice work

OpenStudy (anonymous):

How would I find the slope. And Thanks.

OpenStudy (anonymous):

@satellite73 ?

jimthompson5910 (jim_thompson5910):

the two points are: (12,36) and (18,54) so the slope of the line through these points is m = (y2-y1)/(x2-x1) m = (54-36)/(18-12) m = 18/6 m = 3

OpenStudy (anonymous):

Oh gosh, Thank you so much. So for part 2 I would write it in y = mx+b form?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

So, y=3x, and I don't know what the +b part would be.

jimthompson5910 (jim_thompson5910):

y = mx+b y = 3x + b ... plug in the slope 36 = 3(12) + b ... plug in one of the given points, I'm doing (12,36) solve for b

OpenStudy (anonymous):

And Part 3 is basically solve for b, right?

jimthompson5910 (jim_thompson5910):

no that's still for part 2

OpenStudy (anonymous):

I think I tried solving it wrong. I multiply 3 and 12, right?

jimthompson5910 (jim_thompson5910):

that gives you what

OpenStudy (anonymous):

36

jimthompson5910 (jim_thompson5910):

so 36 = 3(12) + b turns into 36 = 36 + b

OpenStudy (anonymous):

Okay, so I did it correctly.

jimthompson5910 (jim_thompson5910):

what did you get for b

OpenStudy (anonymous):

That's pretty much as far as I got. I don't know what to do next.

jimthompson5910 (jim_thompson5910):

how would you move the 36 on the right side over to the left

OpenStudy (anonymous):

Divide?

jimthompson5910 (jim_thompson5910):

36+b is the same as b+36

jimthompson5910 (jim_thompson5910):

36 = b+36 how do you undo the +36

OpenStudy (anonymous):

Subtract.

jimthompson5910 (jim_thompson5910):

good, giving you b = ???

OpenStudy (anonymous):

0?

jimthompson5910 (jim_thompson5910):

yep, b = 0

jimthompson5910 (jim_thompson5910):

so y = mx+b turns into y = 3x + 0 or just y = 3x

OpenStudy (anonymous):

Okay, how do I do part 3?

jimthompson5910 (jim_thompson5910):

now that you know the equation, you can plug in x = 35 to get y

OpenStudy (anonymous):

y =105?

jimthompson5910 (jim_thompson5910):

good you got it

jimthompson5910 (jim_thompson5910):

so you burn 105 calories when walking 35 min

OpenStudy (anonymous):

Finally. Thanks so much. I can finally go to bed now, hahah.

jimthompson5910 (jim_thompson5910):

lol good night